### 11 Mar 2026 v1

# Distributed Safety Critical Control among Uncontrollable Agents using Reconstructed Control Barrier Functions

Abstract
This paper investigates the distributed safety critical control for multi-agent systems (MASs) in the presence of uncontrollable agents with uncertain behaviors. To ensure system safety, the control barrier function (CBF) is employed in this paper. However, a key challenge is that the CBF constraints are coupled when MASs perform collaborative tasks, which depend on information from multiple agents and impede the design of a fully distributed safe control scheme. To overcome this, a novel reconstructed CBF approach is proposed. In this method, the coupled CBF is reconstructed by leveraging state estimates of other agents obtained from a distributed adaptive observer. Furthermore, a prescribed performance adaptive parameter is designed to modify this reconstruction, ensuring that satisfying the reconstructed CBF constraint is sufficient to meet the original coupled one. Based on the reconstructed CBF, we design a safety-critical quadratic programming (QP) controller and prove that the proposed distributed control scheme rigorously guarantees the safety of the MAS, even in the uncertain dynamic environments involving uncontrollable agents. The effectiveness of the proposed method is illustrated through simulations.
I INTRODUCTION
In recent years, safety critical control has garnered significant attention for its applications in fields such as robotics [1], autonomous driving [2], and aerospace [3]. Beyond established methods such as artificial potential fields (APFs) [4], prescribed performance control (PPC) [5], and nonovershooting control [6], the use of control barrier functions (CBFs) to ensure system safety represents an emerging research direction. The standard CBF-based approach adopt a quadratic program (QP) as the controller to compute the control inputs that strictly satisfies the CBF-induced safety constraint, thereby guaranteeing system safety [7].
Due to the widespread application of multi-agent systems (MAS) in collaborative tasks, ensuring their safety using CBFs has also become an active area of research. The works in [8] and [9] investigate the cooperative control problem for MASs under safety or task constraints using a CBF-based approach. However, both of these studies adopt a centralized control architecture. A distributed CBF-QP control framework is proposed in [10] to address the control problem under collaborative task constraints. Nevertheless, this approach assumes that the MAS for achieving the collaborative task is fully connected. Therefore, designing distributed safety controllers based on CBFs to address the cooperative control problem still remains a challenge.
The primary challenge in designing distributed CBF-based controllers is that the CBF often involves the states of multiple agents, resulting in a coupled CBF. When a standard QP controller is adopted, this coupled CBF induces a coupled constraint that involves the states and control inputs of multiple agents. Consequently, it becomes infeasible for an individual agent to compute a control input that satisfies this constraint using only its local information.
One approach to overcome the issue of coupled CBFs and enable distributed control is the constraint decomposition approach proposed in [11, 12]. The core assumption therein is that a coupled constraint can be decomposed into multiple local constraints, each relying solely on information available to the respective agent. Consequently, if the control input of every agent satisfies its respective local constraint, it can be guaranteed that the original coupled constraint is also satisfied. While the constraint decomposition approach allows for the design of fully distributed QP controllers, a drawback is that not all coupled CBF constraints can be transformed into local ones through a simple additive split as used in [11, 12].
Another challenge arises from the presence of uncontrollable agents with uncertain behavior in the operating environment of MASs [13, 14]. Prominent examples include pedestrians in multi-robot navigation scenarios and human-driven vehicles in autonomous driving contexts. When uncontrollable agents are present in the environment, the constraint decomposition approach proposed in [11, 12] becomes inapplicable. This is because the method requires all agents involved in a coupled constraint to satisfy their respective decomposed local constraints. However, we can only design control inputs for controllable agents, whereas the inputs of uncontrollable agents are entirely unknown and cannot be prescribed. This makes it impossible to satisfy the local constraints corresponding to the uncontrollable agents, which in turn renders the design of distributed safety controllers more challenging.
Contributions: Motivated by the aforementioned limitations, we propose a novel reconstructed CBF approach to resolve the challenge of coupled CBF constraints in distributed safety critical control.
  • By leveraging state estimates from distributed adaptive observers, this method transforms coupled, global CBFs into local ones. We further modify the reconstructed CBFs with a prescribed performance adaptive parameter, which guarantees that satisfying the reconstructed CBF constraints is sufficient to satisfy the original global one.
  • Prior works, such as [11, 12], require the presupposition that all agents involved in a coupled constraint are controllable. In contrast, the reconstructed CBFs method enables controllable agents to compensate for the uncertain behaviors of uncontrollable ones. This relaxes the requirement for global constraint satisfaction, such that instead of requiring all agents to enforce the coupled CBF constraint, only a subset of controllable agents needs to satisfy their reconstructed local constraints.
The remainder of this paper is organized as follows: Section II provides the preliminaries and describes the problem. In Section III, a distributed safety critical control framework based on reconstructed CBFs is proposed. Section IV demonstrates the effectiveness of the proposed method through a simulation example. A conclusion is drawn in Section V.
Notation: Symbols $\mathbb{R}$ represents the real number set. $\mathbb{R}^{n}$ denotes the $n$-dimensional real vector space and $\mathbb{R}^{m\times n}$ denotes the $m\times n$-dimensional real matrix space. $I$ represents an identity matrix with an appropriate dimension. Inequalities $X\succ 0$ and $Y\succeq 0$ indicate that matrices $X$ and $Y$ are positive definite and positive semi-definite, respectively. $\mathbf{1}$ and $\mathbf{0}$ refer to vectors of proper dimensions with all entries to be 0 or $1$, respectively. The Kronecker product of matrices $M\in\mathbb{R}^{m\times n}$ and $N\in\mathbb{R}^{p\times q}$ is denoted by $M\otimes N\in\mathbb{R}^{mp\times nq}$.
II PRELIMINARIES AND PROBLEM STATEMENT
II-A Graph Theory
In this paper, different controllable agents are connected through an undirected graph $\mathcal{G}=\left\{\mathcal{N},\mathcal{E}\right\}$, which consists of agents $\mathcal{N}=\{1,2,\dots,N\}$ and edges $\mathcal{E}=\{(i,j)|i,j\in\mathcal{N}\}$. Since $\mathcal{G}$ is undirected, if the edge $(i,j)\in\mathcal{E}$, then $(j,i)\in\mathcal{E}$. Moreover, $(i,j)\in\mathcal{E}$ indicates that agent $j$ can obtain the state information of agent $i$. In this case, agent $j$ is referred to as a neighbor of agent $i$, and vice versa. Self edge is disallowed in this study, i.e., $(i,i)\notin\mathcal{E},\forall i\in\mathcal{N}$. The adjacency matrix $\mathcal{A}=[a_{ij}]\in\mathbb{R}^{N\times N}$ is given by $a_{ij}=1$ if $(i,j)\in\mathcal{E}$, and $a_{ij}=0$, otherwise. The degree matrix is denoted by $\mathcal{D}=\text{diag}\left\{d_{1},\dots,d_{N}\right\}$, where $d_{i}=\sum_{j=1}^{N}a_{ij}$. The Laplacian matrix $\mathcal{L}$ is provided by $\mathcal{L}=\mathcal{D}-\mathcal{A}$. The undirected graph $\mathcal{G}$ is connected if there is a sequence of connected edges between any agents $i$ and $j$. The set of uncontrollable agents is defined as $\mathcal{V}=\{N+1,\dots,N+V\}$. Similarly, if the controllable agent $i\in\mathcal{N}$ is directly connected with the uncontrollable agent $l\in\mathcal{V}$, which means that agent $i$ is aware of the state information of agent $l$, we have $a_{il}=1$ and $a_{il}=0$, otherwise. Let $\mathcal{N}^{+}:=\mathcal{N}\cup\mathcal{V}$ be the set that includes all controllable agents and uncontrollable agents. For any $i\in\mathcal{N}$ and $j\in\mathcal{N}^{+}$, define $\mathcal{H}_{j}=\mathcal{L}+\mathcal{B}_{j}$ with $\mathcal{B}_{j}=\text{diag}\left\{b_{1j},\dots,b_{Nj}\right\}$, where $b_{ij}=a_{ij}$ if $i\neq j$ and $b_{ii}=1$ otherwise.
II-B System Dynamics and Control Barrier Functions
Consider a class of nonlinear MASs as
$\dot{x}_{i}=f(x_{i})+g(x_{i})u_{i},\,i\in\mathcal{N}^{+},$ (1)
where $x_{i}=\left[x_{i,1},\dots,x_{i,n}\right]^{T}\in\mathbb{R}^{n}$ and $u_{i}\in\mathbb{R}^{m}$ are the state and control input of agent $i$, respectively. The function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is globally Lipschitz continuous with an unknown Lipschitz constant $L_{1}$, while the function $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times m}$ is locally Lipschitz continuous. Let $x=[x_{1}^{T},\dots,x_{N+V}^{T}]^{T}$ and $u=[u_{1}^{T},\dots,u_{N+V}^{T}]^{T}$ denote the stacked state vector and stacked control input vector of all agents in $\mathcal{N}^{+}$, respectively. Note that the control input $u_{l}$ of the uncontrollable agent $l\in\mathcal{V}\subseteq\mathcal{N}^{+}$ is unknown and cannot be designed.
In the CBF-based standard safety critical control method, the fundamental objective is to constrain the system state within a predefined safe set. Given the MAS (1) and a continuously differentiable function $h(x):\mathbb{R}^{(N+V)n}\rightarrow\mathbb{R}$, the safe set $\mathcal{C}$ for (1) is defined as the zero superlevel set of function $h$ as follows:
$\mathcal{C}:=\{x\in\mathbb{R}^{(N+V)n}|h(x)\geq 0\}.$ (2)
The interior and boundary of $\mathcal{C}$, respectively denoted by $\mathrm{Int}(\mathcal{C})$ and $\partial\mathcal{C}$, are defined as follows:
$\mathrm{Int}(\mathcal{C}):=\{x\in\mathbb{R}^{(N+V)n}|h(x)>0\},$
$\partial\mathcal{C}:=\{x\in\mathbb{R}^{(N+V)n}|h(x)=0\}.$ (3)
Afterwards, the definition of CBFs is introduced as follows:
Definition 1 (Zeroing CBFs, ZCBFs [7])
The continuously differentiable function $h(x)$ is a ZCBF if there exists an extended class $\mathcal{K}$ function $\alpha(\cdot)$ for all $x(t)\in\mathcal{C}$, such that
$\sup_{u\in\mathbb{R}^{(N+V)m}}$ $\left[\sum_{i=1}^{N+V}\frac{\partial h}{\partial x_{i}}\left(f(x_{i})+g(x_{i})u_{i}\right)\right]\geq-\alpha(h(x)).$ (4)
Then, the following lemma is established for ZCBFs.
Lemma 1 ([7] )
For a given ZCBF $h(x)$ and MAS (1), if the control input satisfy $u\in\mathcal{U}_{Z}$, where $\mathcal{U}_{Z}$ is the set of control input satisfying (4), the safe set $\mathcal{C}$ is forward invariant.
II-C Problem Statement
In research related to CBFs, QP is commonly employed to synthesize a controller that ensures the control input strictly satisfies safety constraints. For system (1), to guarantee that constraint (4) is strictly satisfied, the standard QP-based controller is formulated as follows:
$u=\mathop{\arg\min}\limits_{{\nu}\in\mathbb{R}^{(N+V)m}}\,\frac{1}{2}(\nu-u_{\text{nom}})^{T}W(\nu-u_{\text{nom}})$ (5a)
$\text{s.t.}\quad\sum_{i=1}^{N+V}\frac{\partial h}{\partial x_{i}}\left(f(x_{i})+g(x_{i})u_{i}\right)\geq-\alpha(h(x)).$ (5b)
The controller (5), also referred to as the safety filter, is used to modify the nominal control input $u_{\text{nom}}\in\mathbb{R}^{(N+V)m}$, which is designed to achieve a primary control task (e.g. trajectory tracking) without considering safety constraints. The remaining terms, $\nu$ and $W\in\mathbb{R}^{(N+V)m\times(N+V)m}$, are the optimization variable and a positive definite weighted matrix, respectively. According to Lemma 1, (5) is able to strictly ensure the safety of the system while achieving the control task corresponding to $u_{\text{nom}}$ as effectively as possible. However, the presence of the coupled constraint (4) renders the controller (5) centralized. Besides, it is not possible to solve for the control input that satisfies (4) using only the information from a single agent. Moreover, due to the presence of uncontrollable agents, $u_{i}$ for $i\in\mathcal{V}$ cannot be designed, which renders (5) inapplicable and exacerbates the difficulty in designing a distributed safety controller.
To overcome the aforementioned challenges, the control objective of this paper is as follows:
  • Design a distributed QP controller for each controllable agent $i\in\mathcal{N}$, which rigorously guarantees the safety constraints while utilizing only the local information of the agent;
  • The designed distributed safety controller remains applicable even when the safety constraints involve uncontrollable agents, i.e., $\frac{\partial h}{\partial x_{i}}\not\equiv\boldsymbol{0}^{T}$ for $i\in\mathcal{V}$.
Remark 1
Although our analysis, as presented in (5b), focuses on the case where all controllable agents are subject to a single, common coupled CBF constraint, the proposed method can be readily extended to scenarios involving multiple heterogeneous coupled CBF constraints among agents. We demonstrate this capability with a case study in Section IV.
To achieve the control objectives, the following assumptions are imposed.
Assumption 1
The undirected graph $\mathcal{G}$ is connected, and the state of uncontrollable agent $l$ for any $l\in\mathcal{V}$ can be obtained by at least one controllable agent.
Assumption 2
The control effectiveness $g(x_{i})u_{i}$ is bounded, i.e., $\left\lVert g(x_{i})u_{i}\right\rVert\leq D$ for any $i\in\mathcal{N}^{+}$, where $D>0$ is an unknown upper bound.
Assumption 3
The derivative of the function $h(z)$, i.e., $\frac{dh(z)}{dz}$, is designed to be locally Lipschitz continuous.
Assumption 4
For the CBF $h(x)$, both its function form and its value are known to any controllable agent $i\in\mathcal{N}$.
Assumption 5
At the initial time $t=0$, each agent $i$ is in the interior of the safe set $\mathcal{C}$, i.e., $h(x(0))>0$.
Remark 2
Assumption 1 relaxes the fully connected topology required in [10] while maintaining the necessary connectivity for cooperative tasks like consensus [15]. Assumption 2 standardly bounds control effort [16, 17], and Assumption 3 is easily satisfied. This work departs from [11] which assumes that coupled constraints can be decomposed into local constraints known to individual agents. Instead, Assumption 4 posits that an agent must be aware of both the safety objective (the functional form of $h$) and its current safety status (the value of $h$). This is reasonable. For instance, in collision avoidance ($h(x_{1},x_{2}):=\lVert x_{1}-x_{2}\rVert^{2}-d^{2}$, where $d>0$ is a safe distance), agents do not need each other’s full state; the CBF value (relative distances) obtained via local sensing (e.g., ultra-wideband) are sufficient. Finally, Assumption 5 necessitates initial safety to prevent immediate constraint violations caused by uncertainties [2].
The following lemmas will be utilized in the subsequent sections.
Lemma 2 ([18])
Under Assumption 1, it is clearly that the non-zero diagonal matrix $\mathcal{B}_{j}\succeq 0$ for all $j\in\mathcal{N}^{+}$. Thus, it can be obtained that $\mathcal{H}_{j}\succ 0$, with the eigenvalues $\lambda_{j,l}$ of $\mathcal{H}_{j}$ satisfying $0<\lambda_{j,1}\leq\dots\leq\lambda_{j,N}$.
Lemma 3 ([19])
The following inequality holds for any $\epsilon>0$ and any $\varpi\in\mathbb{R}$:
$|\varpi|\leq\frac{\varpi^{2}}{\sqrt{\varpi^{2}+\epsilon^{2}}}+\epsilon.$ (6)
III MAIN RESULTS
The distributed control scheme proposed in this paper will be presented in this section.
III-A Design of Distributed Adaptive Observer
Noting that constraint (5b) incorporates the states and control inputs of multiple agents, the implementation of distributed control becomes challenging due to the incomplete information acquisition by a single agent. To overcome this issue, we first design a distributed observer for each agent $i\in\mathcal{N}$, which allows it to estimate the states of the other agents.
$\dot{\hat{x}}_{i,l}=f(\hat{x}_{i,l})-\hat{\delta}_{i,l}\xi_{i,l},$
$\xi_{i,l}=\sum_{j=1}^{N}a_{ij}\left(\hat{x}_{i,l}-\hat{x}_{j,l}\right)+b_{il}\left(\hat{x}_{i,l}-x_{l}\right),$
$\dot{\hat{\delta}}_{i,l}=2\xi_{i,l}^{T}P_{l}\xi_{i,l}-\sigma_{l}{\hat{\delta}}_{i,l},\,\,l=1,\dots,N+V,$ (7)
where $\hat{x}_{i,l}$ is the estimate of the state $x_{l}$ by agent $i$, $P_{l}\in\mathbb{R}^{n\times n}$ is a symmetric positive definite matrix, $\sigma_{l}$ is a positive constant. $\hat{\delta}_{i,l}$ is an adaptive parameter used to estimate the constant $\delta_{l}$, which satisfies
$2\lambda_{l,N}P_{l}^{2}+\frac{L_{1}^{2}}{\lambda_{l,1}}I-2\lambda_{l,1}\delta_{l}P_{l}+\sigma_{l}P_{l}\prec 0.$ (8)
Afterwards, the following proposition can be established.
Proposition 1
Consider the MAS described by (1) with the distributed adaptive observers (III-A). Provided that Assumptions 1 and 2 are satisfied, it can be concluded that
[…]
2) if the signal $x_{l}$ for $l\in\mathcal{N}^{+}$ is bounded, then all signals of the observer (III-A) are bounded.
Proof:
The time derivative of $\tilde{x}_{i,l}$ is calculated as
$\dot{\tilde{x}}_{i,l}=\tilde{f}_{i,l}-d_{l}-\hat{\delta}_{i,l}\xi_{i,l},$ (9)
where $\tilde{f}_{i,l}:=f(\hat{x}_{i,l})-f(x_{i,l})$ and $d_{l}:=g({x}_{l})u_{l}$. Consequently, define $\tilde{x}_{l}:=\left[\tilde{x}_{1,l}^{T},\dots,\tilde{x}_{N,l}^{T}\right]^{T}$. Consider the following Lyapunov function
$V_{l}=\tilde{x}_{l}^{T}\left(\mathcal{H}_{l}\otimes P_{l}\right)\tilde{x}_{l}+\frac{1}{2}\sum_{i=1}^{N}\tilde{\delta}_{i,l}^{2},$ (10)
where $\tilde{\delta}_{i,l}=\delta_{l}-\hat{\delta}_{i,l}$. It can be derived that
$\dot{\tilde{x}}_{l}=\tilde{F}_{l}-\boldsymbol{1}\otimes d_{l}-\hat{\delta}_{l}\left(\mathcal{H}_{l}\otimes I\right)\tilde{x}_{l},$ (11)
where $\tilde{F}_{l}=\left[\tilde{f}_{1,l}^{T},\dots,\tilde{f}_{N,l}^{T}\right]^{T}$ and $\hat{\delta}_{l}=\mathrm{diag}\{\underbrace{\hat{\delta}_{1,l},\dots,\hat{\delta}_{1,l},}\limits_{n}\underbrace{\hat{\delta}_{2,l},\dots,\hat{\delta}_{2,l},}\limits_{n}\dots\underbrace{,\hat{\delta}_{N,l},\dots,\hat{\delta}_{N,l}}\limits_{n}\}_{Nn\times Nn}$. Thus, the derivative of $V_{l}$ along (11) is obtained as
$\dot{V}_{l}=$ $2\tilde{x}_{l}^{T}\left(\mathcal{H}_{l}\otimes P_{l}\right)\tilde{F}_{l}-2\tilde{x}_{l}^{T}\left(\mathcal{H}_{l}\otimes P_{l}\right)\left(\boldsymbol{1}\otimes d_{l}\right)$
$-2\tilde{x}_{l}^{T}\left(\mathcal{H}_{l}\otimes P_{l}\right)\hat{\delta}_{l}\left(\mathcal{H}_{l}\otimes I\right)\tilde{x}_{l}-\sum_{i=1}^{N}\tilde{\delta}_{i,l}\dot{\hat{\delta}}_{i,l}$
[…]
$-2\sum_{i=1}^{N}\hat{\delta}_{i,l}\xi_{i,l}^{T}P_{l}\xi_{i,l}-\sum_{i=1}^{N}\tilde{\delta}_{i,l}\dot{\hat{\delta}}_{i,l}.$ (12)
Based on the global Lipschitz continuity of the function $f(\cdot)$ and Assumption 4, it can be deduced that
[…]
Substituting (13) into (III-A) and using Lemma 2 yields:
$\dot{V}_{l}\leq$ $2\tilde{x}_{l}^{T}\left(\mathcal{H}_{l}^{2}\otimes P_{l}^{2}\right)\tilde{x}_{l}+L_{1}^{2}\tilde{x}_{l}^{T}\tilde{x}_{l}+ND^{2}$
$-2\sum_{i=1}^{N}\left(\delta_{l}-\tilde{\delta}_{i,l}\right)\xi_{i,l}^{T}P_{l}\xi_{i,l}-\sum_{i=1}^{N}\tilde{\delta}_{i,l}\dot{\hat{\delta}}_{i,l}$
$\leq$ $2\lambda_{l,N}\tilde{x}_{l}^{T}\left(\mathcal{H}_{l}\otimes P_{l}^{2}\right)\tilde{x}_{l}+\frac{L_{1}^{2}}{\lambda_{l,1}}\tilde{x}_{l}^{T}\left(\mathcal{H}_{l}\otimes I\right)\tilde{x}_{l}$
$-2\lambda_{l,1}\delta_{l}\tilde{x}_{l}^{T}\left(\mathcal{H}_{l}\otimes P_{l}\right)\tilde{x}_{l}+\sigma_{l}\sum_{i=1}^{N}\tilde{\delta}_{i,l}\hat{\delta}_{i,l}+ND^{2}.$ (14)
Combining (8) and (III-A), it is deduced that
$\dot{V}_{l}\leq$ $-\sigma_{l}\tilde{x}_{l}^{T}\left(\mathcal{H}_{l}\otimes P_{l}\right)\tilde{x}_{l}-\frac{\sigma_{l}}{2}\sum_{i=1}^{N}\tilde{\delta}_{i,l}^{2}+\frac{\sigma_{l}}{2}\sum_{i=1}^{N}\delta_{l}^{2}+ND^{2}$
$=-\sigma_{l}V_{l}+\Xi_{l}.$ (15)
Performing integration on both sides of (III-A) yields that
$V_{l}(t)\leq\left(V_{l}(0)-\frac{\Xi_{l}}{\sigma_{l}}\right)\mathrm{e}^{-\sigma_{l}t}+\frac{\Xi_{l}}{\sigma_{l}}.$ (16)
From (10) and (16), it can be concluded that $V_{l}$, $\tilde{x}_{i,l}$ and $\tilde{\delta}_{i,l}$ for $i\in\mathcal{N},l\in\mathcal{N}^{+}$ remain bounded. Moreover, $\hat{\delta}_{i,l}$ and $\xi_{i,l}$ are also bounded. From (11) and (13), all $\dot{\tilde{x}}_{i,l}$ remain bounded. According to (16), it also can be obtain that
[…]
Furthermore, if $x_{l}$ is bounded, then, based on the previous analysis, it is evident that all the signals of the observer (III-A) are bounded. ∎
III-B Reconstructed Control Barrier Functions
Based on the state estimate from the observer (III-A), the coupled CBF $h(x)$ is then reconstructed for each agent $i\in\mathcal{N}$ as follows:
$\hat{h}_{i}(\hat{\upsilon}_{i},\vartheta_{i})=h(\hat{\upsilon}_{i})-\vartheta_{i},$
$\dot{\vartheta}_{i}=\begin{aligned} &-c_{i}\frac{e_{i}(\rho_{i}-e_{i})}{\rho_{i}}\varepsilon_{i}-\frac{\varrho e_{i}}{\rho_{i}}(\rho_{0}^{i}-\rho_{\infty}^{i})\mathrm{e}^{-\varrho t}\\ &-\frac{\rho_{i}\varepsilon_{i}}{4e_{i}(\rho_{i}-e_{i})}-\frac{\hat{r}_{i}^{2}\left(\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert+\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)\chi_{i}}{\sqrt{\chi_{i}^{2}\hat{r}_{i}^{2}+\epsilon_{i}^{2}}},\end{aligned}$
$\dot{\hat{r}}_{i}=\frac{\gamma_{i}|\varepsilon_{i}|\rho_{i}}{2e_{i}(\rho_{i}-e_{i})}\left(\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert+\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)-\varsigma_{i}\hat{r}_{i},$ (18)
where $\hat{\upsilon}_{i}$ is the vector obtained by replacing the element $x_{j}$ ($j\neq i,j\in\mathcal{N}^{+}$) in the vector $x$ with $\hat{x}_{i,j}$, $\hat{\bar{x}}_{i}:=\left[\hat{x}_{i,1}^{T},\dots,\hat{x}_{i,N+V}^{T}\right]^{T}$ and $\chi_{i}:=\frac{\varepsilon_{i}\rho_{i}}{2e_{i}(\rho_{i}-e_{i})}\left(\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert+\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)$. The parameters $c_{i}$, $\varrho$, $\epsilon_{i}$, $\varsigma_{i}$ and $\gamma_{i}$ are positive constants. Moreover, $\hat{h}_{i}$ is the reconstructed CBF, $\vartheta_{i}$ and $\hat{r}_{i}$ are adaptive parameters, $e_{i}$ is the reconstruction error, given by
$e_{i}:=h(x)-\hat{h}_{i}(\hat{\upsilon}_{i},\vartheta_{i}),$ (19)
and $\varepsilon_{i}$ and $\rho_{i}$ are respectively defined as
$\varepsilon_{i}=\frac{1}{2}\ln\left(\frac{e_{i}}{\rho_{i}-e_{i}}\right),\,\rho_{i}(t)=\left(\rho_{0}^{i}-\rho_{\infty}^{i}\right)\text{e}^{-\varrho t}+\rho_{\infty}^{i}.$
It should be noted that $\vartheta_{i}$ is an adaptive parameter designed based on PPC theory. Its role is to modify the function $h(\hat{\upsilon}_{i})$ in (III-B) to ensure that when the reconstructed CBF constraint $\hat{h}_{i}\geq 0$ is satisfied, the original coupled CBF constraint $h\geq 0$ also holds. This property will be formally proven in the subsequent theorem. The initial value of $\vartheta_{i}$, $\rho_{0}^{i}$ and $\rho_{\infty}^{i}$ are chosen such that $0<\rho_{\infty}^{i}<\rho_{0}^{i}$, $0<e_{i}(0)<\rho_{0}^{i}$ and $\hat{h}_{i}(\hat{\upsilon}_{i}(0),\vartheta_{i}(0))\geq 0$.
Then, regarding the reconstructed CBF (III-B), the following result can be established.
Theorem 1
Considering reconstructed CBF (III-B), under Assumption 1-5, it can be concluded that:
1) The reconstruction error $e_{i}$ (19) satisfies the prescribed performance constraint, i.e., $0<e_{i}(t)<\rho_{i}(t)$;
2) If $\hat{h}_{i}\geq 0$ holds for at least one $i\in\mathcal{N}$, then it follows that $h(x)>0$.
Proof:
1) Using proof by contradiction, we assume that $e_{i}$ first exceeds the performance bounds $(0,\rho_{i})$ at some finite time $\bar{t}$, which indicates that $\lim_{t\to\bar{t}}\varepsilon_{i}(t)\rightarrow\infty$. Hence, for $t\in[0,\bar{t})$, we have $e_{i}\in(0,\rho_{i})$. The following Lyapunov function is selected as
$V_{i}=\frac{1}{2}\varepsilon_{i}^{2}+\frac{1}{2\gamma_{i}}\tilde{r}_{i}^{2},$ (20)
where $\tilde{r}_{i}=r_{i}-\hat{r}_{i}$. Positive constant $r_{i}$ satisfies a certain condition to be introduced later. Taking the derivative of $V_{i}$ over the interval $0,\bar{t})$ yields
$\dot{V}_{i}=$ $\varepsilon_{i}\left[\frac{-1}{2\left(\rho_{i}-e_{i}\right)}\dot{\rho}_{i}+\frac{\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\right$.
$\left.\times\left(\dot{\vartheta}_{i}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\dot{\hat{\upsilon}}_{i}+\frac{dh(x)}{dx}\dot{x}\right)\right]-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}$
$=$ $\varepsilon_{i}\left(\frac{\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\dot{\vartheta}_{i}-\frac{1}{2\left(\rho_{i}-e_{i}\right)}\dot{\rho}_{i}\right)$
$+\frac{\varepsilon_{i}\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left[\frac{dh(x)}{dx}\left(\dot{x}_{i}-\dot{\hat{\upsilon}}_{i}\right)\right$.
$\left.+\left(\frac{dh(x)}{dx}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right)\dot{\hat{\upsilon}}_{i}\right]-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}$. (21)
Combining Proposition 1 and Assumption 3, it is deduced that
[…]
where $L_{2}$ is a Lipschitz constant of $\frac{dh(z)}{dz}$, $\Delta_{i,1}$ and $\Delta_{i,2}$ are positive constants. Defining $r_{i}:=\max\{\Delta_{i,1},\Delta_{i,2}\}$ and applying Lemma 3, it can be obtained that
$\dot{V}_{i}\leq$ $\varepsilon_{i}\left(\frac{\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\dot{\vartheta}_{i}-\frac{1}{2\left(\rho_{i}-e_{i}\right)}\dot{\rho}_{i}\right)$
$+\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left[\left(\left\lVert\frac{dh(x)}{dx}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}\right\rVert\right)r_{i}\right]-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}$
$\leq$ $\frac{\rho_{i}\varepsilon_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\dot{\vartheta}_{i}-\frac{\varepsilon_{i}}{2\left(\rho_{i}-e_{i}\right)}\dot{\rho}_{i}+\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}(\rho_{i}-e_{i})}$
$\times\left[\left(\left\lVert\frac{dh(x)}{dx}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}\right\rVert-\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert-\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}\right]$
$+\frac{|\varepsilon_{i}|\rho_{i}\tilde{r}_{i}}{2e_{i}(\rho_{i}-e_{i})}\left(\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert+\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)$
$+\frac{\chi_{i}^{2}\hat{r}_{i}^{2}}{\sqrt{\chi_{i}^{2}\hat{r}_{i}^{2}+\epsilon_{i}^{2}}}+\epsilon_{i}-\frac{1}{\gamma_{i}}\tilde{r}_{i}\dot{\hat{r}}$. (24)
Applying the reverse triangle inequality yields that
$\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left[\left(\left\lVert\frac{dh(x)}{dx}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}\right\rVert-\left\lVert\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert-\left\lVert\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}\right]$
$\leq\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\left[\left(\left\lVert\frac{dh(x)}{dx}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}-\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}\right]$
$\leq\frac{1}{2}\left(\frac{|\varepsilon_{i}|\rho_{i}}{2e_{i}\left(\rho_{i}-e_{i}\right)}\right)^{2}+\frac{1}{2}\Delta_{i}^{2},$ (25)
where $\Delta_{i}\geq\left(\left\lVert\frac{dh(x)}{dx}-\frac{dh(\hat{\upsilon}_{i})}{d\hat{\upsilon}_{i}}\right\rVert+\left\lVert\dot{\hat{\upsilon}}_{i}-\dot{\hat{\bar{x}}}_{i}\right\rVert\right)r_{i}$ is a positive constant. Substituting (III-B), (III-B) into (III-B) and using Young inequality yields that
$\dot{V}_{i}\leq$ $-\frac{c_{i}}{2}\varepsilon_{i}^{2}+\varsigma_{i}\tilde{r}_{i}\hat{r}_{i}+\frac{1}{2}\Delta_{i}^{2}+\epsilon_{i}$
$\leq$ $-\frac{c_{i}}{2}\varepsilon_{i}^{2}-\frac{\varsigma_{i}}{2}\tilde{r}_{i}^{2}+\frac{\varsigma_{i}}{2}{r}_{i}^{2}+\frac{1}{2}\Delta_{i}^{2}+\epsilon_{i}$
$\leq$ $-\zeta_{i}V_{i}+\Omega_{i},$ (26)
where $\zeta_{i}=\min\{c_{i},\varsigma_{i}\gamma_{i}\}$ and $\Omega_{i}:=\frac{\varsigma_{i}}{2}{r}_{i}^{2}+\frac{1}{2}\Delta_{i}^{2}+\epsilon_{i}$ is a positive constant. Integrating both sides of (III-B), it is obtained that
$V_{i}(t)\leq\left(V_{i}(0)-\frac{\Omega_{i}}{\zeta_{i}}\right)\mathrm{e}^{-\zeta_{i}t}+\frac{\Omega_{i}}{\zeta_{i}}.$ (27)
From (27), it can be known that $V_{i}(\bar{t}^{-})$ is bounded, i.e., $\varepsilon_{i}(\bar{t}^{-})$ is bounded, which is in contradiction with the assumption $\lim_{t\to\bar{t}}\varepsilon_{i}(t)\rightarrow\infty$. Therefore, for $t\in[0,+\infty)$, $\varepsilon_{i}$ and $\tilde{r}_{i}$ remain bounded, i.e., reconstruction error $e_{i}$ satisfy the prescribed performance criteria.
2) Since for $t\in[0,+\infty)$, we have $e_{i}>0$, which implies that $h(x)>\hat{h}_{i}(\hat{\upsilon}_{i},\vartheta_{i})$ for any $i\in\mathcal{N}$ holds. Thus, if $\hat{h}_{i}(\hat{\upsilon}_{i},\vartheta_{i})\geq 0$ holds for at least one $i\in\mathcal{N}$, it is deduced that $h(x)>0$ holds. ∎
Theorem 1.2) stems from the fact clarified in Remark 1: all $\hat{h}_{i}$, $i\in\mathcal{N}$, are reconstructions of the original coupled CBF $h$. Moreover, for scenarios with multiple heterogeneous coupled CBFs, the non-negativity of each original CBF would need to be individually guaranteed by its respective reconstructed CBF remaining non-negative, which can be easily extended from our approach by having each agent reconstruct its own original coupled CBF via (III-B).
III-C Design of Distributed Safety Controller
To guarantee the overall safety of MAS (1), based on the distributed observer (III-A) and the reconstructed CBF (III-B), a distributed safety controller for agent $i\in\mathcal{N}$ is designed as follows:
$u_{i}=\mathop{\arg\min}\limits_{{\nu_{i}}\in\mathbb{R}^{m}}\,\frac{1}{2}(\nu_{i}-u_{i,\text{nom}})^{T}W_{i}(\nu_{i}-u_{i,\text{nom}})$ (28a)
$\begin{aligned} \text{s.t.}\quad&\frac{\partial\hat{h}_{i}}{\partial x_{i}}(f(x_{i})+g(x_{i})u_{i})+\sum_{l\in\mathcal{N}^{+},l\neq i}\frac{\partial\hat{h}_{i}}{\partial\hat{x}_{i,l}}\dot{\hat{x}}_{i,l}-\dot{\vartheta}_{i}\\ &\geq-\alpha_{i}(\hat{h}_{i}),\end{aligned}$ (28b)
where $\nu_{i}$ represents the optimization variable, $u_{i,\text{nom}}$ is the nominal control input of agent $i$, $W_{i}\in\mathbb{R}^{m\times m}$ is a positive definite weighted matrix, and $\alpha_{i}(\cdot)$ is an extended class $\mathcal{K}$ function. Then, we have the following theorem.
Theorem 2
Considering the MAS modeled by (1) under Assumption 1-5. With the proposed distributed control scheme (III-A), (III-B) and (28), the safe set $\mathcal{C}$ is forward invariant, i.e., the safety of all agents in $\mathcal{N}$ is guaranteed.
Proof:
At initial time $t=0$, according to Assumption 5, we have $h(x(0))>0$. By appropriately selecting the parameters, we can always ensure that $\hat{h}_{i}(\hat{\upsilon}_{i}(0),\vartheta_{i}(0))\geq 0$ and $0<e_{i}(0)<\rho_{0}^{i}$. As a concrete example, we can set $\rho_{0}^{i}=h(x(0))$ and $\vartheta_{i}(0)=h(\hat{\upsilon}_{i}(0))-h(x(0))/2$, then, it follows that $0<e_{i}(0)=h(x(0))/2<h(x(0))=\rho_{0}^{i}$ and $\hat{h}_{i}(\hat{\upsilon}_{i}(0),\vartheta_{i}(0))=h(x(0))/2\geq 0$. Furthermore, if $\hat{h}_{i}(\hat{\upsilon}_{i}(0),\vartheta_{i}(0))\geq 0$, from Lemma 1, it can be known that $\hat{h}_{i}(\hat{\upsilon}_{i},\vartheta_{i})\geq 0$ always holds for $t\in[0,+\infty)$ with the controller (28). Recalling Theorem 1, it can be inferred that $h(x(t))\geq 0$ also holds for $t\in[0,+\infty)$, which implies that the safety set $\mathcal{C}$ is forward invariant. Therefore, the safety of all agents in $\mathcal{N}$ is rigorously guaranteed. ∎
Remark 3
The proposed fully distributed scheme rigorously guarantees safety by rendering $\mathcal{C}$ forward invariant (Theorem 2). Crucially, unlike methods [11, 12] that require all agents in a coupled constraint to actively comply, our reconstructed CBF approach accommodates uncontrollable agents. As shown in Theorem 1, maintaining the reconstructed constraint (e.g., $\hat{h}_{i}\geq 0$) ensures the original coupled constraint $h\geq 0$ is satisfied. Therefore, controllable agents can compensate for uncertain behaviors of uncontrollable agents, requiring only a subset of agents to actively satisfy the constraints.
IV SIMULATION
In this section, the effectiveness of the proposed method is demonstrated by a multi-robot navigation case study. Consider an MAS consisting of four robots with $\mathcal{N}^{+}=\{1,2,3,4\}$, where robot $4$ is an uncontrollable agent, while the remaining agents are controllable. The dynamics of robots are given as
$\begin{cases}\dot{\boldsymbol{p}}_{i}=\left[\begin{matrix}\cos\theta_{i}&-l\sin\theta_{i}\\ \sin\theta_{i}&l\cos\theta_{i}\end{matrix}\right]u_{i},\\ \dot{\theta}_{i}=\omega_{i},\end{cases}i=1,2,3,4,$ (29)
where ${\boldsymbol{p}}_{i}=[x_{i},y_{i}]^{T}$ denotes the position of the center of robot $i$ and $\theta_{i}$ is its heading angle. $u_{i}=[v_{i},\omega_{i}]^{T}$ is the control input of robot $i$, where $v_{i}$ and $\omega_{i}$ are the linear velocity and angular velocity of robot $i$, respectively. $l=0.036$ is the distance between the center point ${\boldsymbol{p}}_{i}$ and the wheel axle of robot $i$. As the subsequent robotic task depends only on ${\boldsymbol{p}}_{i}$, (29) can be regarded as (1) with $f=\boldsymbol{0}$ and $g=[\cos\theta_{i},-l\sin\theta_{i};\sin\theta_{i},l\cos\theta_{i}]$. The communication topology is illustrated in Fig. 1.
Figure 1: Communication topology.
The robots’ working environment contains three circular obstacles $O_{i}$, $i=1,\dots,3$. The center position ${\boldsymbol{p}}_{O,i}$ and the radius $r_{O,i}$ of $O_{i}$ is set as: ${\boldsymbol{p}}_{O,1}^{T},r_{O,1}]=[0.8,2.5,0.5]$, ${\boldsymbol{p}}_{O,2}^{T},r_{O,2}]=[3,3.5,0.3]$, ${\boldsymbol{p}}_{O,3}^{T},r_{O,3}]=[3,1.5,0.5]$. Besides, the environment includes a goal point $G_{0}$ located at $\boldsymbol{p}_{G,0}=[0,4]^{T}$ and a circular target region $G_{1}$ centered at $\boldsymbol{p}_{G,1}=[1,0.5]^{T}$ with a radius $r_{G,1}=0.4$. Robot $4$ is required to navigate to the goal point $G_{0}$ and robots $1,2$ are tasked with navigating to the goal region $G_{1}$. In addition, robots $i$ for $i=1,2,3$ are subject to the following constraints:
  • Collision avoidance: Robots $i\in\{1,2,3\}$ must avoid obstacles and other robots, yielding the constraints:
[…]
    where $r_{R}=0.1$ is the collision radius of the robots.
  • Coupled constraints: Robots $1$ and $2$ must cooperate, while robot $3$ follows the uncontrollable robot $4$. Both pairs must maintain a distance within $d_{f}=1.25$, yielding:
[…]
Then, we derive the following CBF:
$h_{i}=-\frac{1}{20}\ln\left(\sum_{l=1}^{3}\text{e}^{-20b_{i,O_{l}}}+\sum_{j\neq i,j\in\mathcal{N}^{+}}\text{e}^{-20b_{i,j}}+\text{e}^{-20b_{i}}\right).$
It is deduced that $h_{i}\leq\min_{l\in\{1,2,3\},j\neq i,j\in\mathcal{N}^{+}}\{b_{i,O_{l}},b_{i,j},b_{i}\}$. In the considered case study, there exists multiple heterogeneous coupled CBFs, i.e., $h_{i}$, $i=1,2,3$. Nevertheless, we can still reconstruct $h_{i}$ using (III-B) and subsequently establish a safety controller based on (28).
Figure 2: Trajectories of 4 robots.
Figure 3: Reconstructed CBF $\hat{h}_{i}$ and reconstruction error $e_{i}$, $i=1,\dots,3$.
The initial states of the robots are set as follows: $\boldsymbol{p}_{1}(0)^{T},\theta_{1}(0)]=[4,4,0]$, $\boldsymbol{p}_{2}(0)^{T},\theta_{2}(0)]=[4,3,0]$, $\boldsymbol{p}_{3}(0)^{T},\theta_{3}(0)]=[4,0.5,-\pi]$, $\boldsymbol{p}_{4}(0)^{T},\theta_{4}(0)]=[3.5,0.6,-\pi]$. The distributed observers here are used solely for estimating $\boldsymbol{p}_{i}$ and the distributed observers parameters are set as: the initial estimate $\hat{x}_{i,j}$ is set to be identical to the actual initial state, $\hat{\delta}_{i,j}(0)=2$, $\sigma_{j}=0.01$, $P_{j}=\mathrm{diag}\{2,2\}$ for $i=1,2,3,j\in\mathcal{N}^{+}$. The parameters for the reconstructed CBFs are set as follows: $\theta_{i}(0)=0.1$, $\hat{r}_{i}(0)=0$, $\rho_{0}^{i}=1$, $\rho_{\infty}^{i}=0.15$, $\varrho=1$, $c_{i}=0.01$, $\varsigma_{i}=0.8$, $\gamma_{i}=0.01$, $\epsilon_{i}=0.01$ for $i=1,2,3$. For QP controller, we choose $\alpha_{1}(\hat{h}_{1})=\hat{h}_{1}$, $\alpha_{2}(\hat{h}_{2})=\hat{h}_{2}$, $\alpha_{3}(\hat{h}_{3})=\hat{h}_{3}^{5}/10$, and $W_{i}=g(\theta_{i})^{T}g(\theta_{i})$ for $i=1,2,3$. Additionally, the constraints on the robots’ maximum linear velocity and angular velocity are imposed, such that $|v_{i}|\leq v_{\max}$ and $|\omega_{i}|\leq\omega_{\max}$ for $i\in\mathcal{N}^{+}$, where $v_{\max},\omega_{\max}]=[0.22,2.84]$.
Figure 4: Control input $u_{i}$, $i=1,\dots,4$.
The trajectories of the robots are presented in Fig. 2. Fig. 3 illustrates the reconstructed CBFs $\hat{h}_{i}$ and reconstruction errors $e_{i}$, $i=1,2,3$, where the black dashed line represents the prescribed performance boundary. The control inputs for each robot are displayed in Fig. 4. As shown in Fig. 3, all reconstructed CBFs remain non-negative and the reconstruction error satisfies the prescribed performance requirements. This implies that the original constraint $h_{i}\geq 0$ is also satisfied. Therefore, the proposed distributed control method effectively ensures system safety.
V CONCLUSIONS
This paper presented a distributed safety critical control scheme based on reconstructed CBFs to address the control problem of MASs subject to coupled safety constraints. The original coupled CBF is reconstructed using state estimates obtained from a distributed adaptive observer, which enables the resulting reconstructed CBFs to be fully distributed. The reconstruction process is modified by designing a prescribed performance adaptive parameter, such that the satisfaction of the reconstructed CBF constraints guarantees that of the original coupled one. Based on the reconstructed CBFs, a safety QP controller is designed and we prove that this controller strictly guarantees the safety of the MAS. Compared to previous works, our proposed method is also applicable to uncertain dynamic environments containing uncontrollable agents. A simulation is conducted to demonstrate the effectiveness of the proposed method.